# Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication

### v1

@@ 3.3 Latent space communication
While each GAE independently produces a latent space for its respective graph—resulting in misaligned embeddings—we address this limitation by incorporating deep functional maps to learn explicit mappings between latent representations. Rather than directly comparing raw embeddings—which may differ by arbitrary isometric transformations—we learn functional maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ that transform functions between latent spaces. These maps are optimized within our network using Equation (1), where $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ represent embeddings from the shared dual-pass encoder.
To facilitate latent space communication, our framework leverages spectral geometry principles and a regularized functional map module that enforces structural constraints. Specifically, we impose bijectivity and orthogonality losses to ensure the maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ are approximately invertible and locally isometric, preserving essential geometric properties.
The bijectivity loss promotes invertibility by ensuring that functions mapped from one latent space to the other and back are reconstructed accurately, thereby enforcing structural consistency and encouraging the learned feature spaces to be approximately invertible and mutually aligned through the learned maps. Formally, it is defined as:
$\mathcal{L}_{\text{bij}}=\|\mathbf{C}_{12}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}\mathbf{C}_{12}-\mathbf{I}\|_{F}^{2}.$ (3)
@@
$\mathcal{L}_{\text{orth}}=\|\mathbf{C}_{12}\mathbf{C}_{12}^{\top}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}^{\top}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}.$ (4)
@@ 3.5 Model optimization and training Loss
Our model jointly optimizes GAE parameters and functional maps $\mathbf{C}_{12}$, $\mathbf{C}_{21}$ through end-to-end training. Given embeddings $\mathbf{Z}_{1}=f_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})$ and $\mathbf{Z}_{2}=f_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})$ from the shared graph encoder, we project them into spectral domains using graph Laplacian eigenvectors, yielding descriptors $\hat{\mathbf{F}}_{1}$ and $\hat{\mathbf{F}}_{2}$. The functional maps $\mathbf{C}_{12}\in\mathbb{R}^{k\times k}$ and $\mathbf{C}_{21}\in\mathbb{R}^{k\times k}$ align these spectral features via:
$\mathcal{L}_{\text{FM}}^{12}=\alpha\left\|\mathbf{C}_{12}\hat{\mathbf{F}}_{1}-\hat{\mathbf{F}}_{2}\right\|_{F}^{2}+\beta\left\|\Lambda_{2}\mathbf{C}_{12}-\mathbf{C}_{12}\Lambda_{1}\right\|_{F}^{2}$ (5)
$\mathcal{L}_{\text{FM}}^{21}=\alpha\left\|\mathbf{C}_{21}\hat{\mathbf{F}}_{2}-\hat{\mathbf{F}}_{1}\right\|_{F}^{2}+\beta\left\|\Lambda_{1}\mathbf{C}_{21}-\mathbf{C}_{21}\Lambda_{2}\right\|_{F}^{2}$ (6)
We incorporate these objectives as differentiable loss terms, treating $\mathbf{C}_{12}$ and $\mathbf{C}_{21}$ as trainable parameters that are jointly optimized end-to-end via backpropagation.
These losses are combined with the standard GAE reconstruction loss, which minimizes binary cross-entropy between the observed adjacency matrix $\mathbf{A}$ and its reconstruction $\hat{\mathbf{A}}$,
$\mathcal{L}_{\text{rec}}=\mathrm{BCE}(\mathbf{A}_{s},\hat{\mathbf{A}}_{s})+\mathrm{BCE}(\mathbf{A}_{t},\hat{\mathbf{A}}_{t}),$ (7)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{FM}}\left(\mathcal{L}_{\text{FM}}^{12}+\mathcal{L}_{\text{FM}}^{21}\right)+\lambda_{\text{bij}}\,\mathcal{L}_{\text{bij}}+\lambda_{\text{orth}}\,\mathcal{L}_{\text{otrh}}$ (8)
@@ 4 Experiments
In this section, we aim to address the following research questions: (1) robustness: is GADL more robust to feature and structural inconsistencies compared to existing state-of-the-art graph alignment methods? (2) effectiveness: does GADL outperform state-of-the-art methods on real-world graph alignment tasks? (3) generalization: How effectively can GADL generalize to vision–language alignment task? (4) ablation analysis: what is the contribution of each component in GADL to the overall alignment performance? (5) encoder evaluation: how does the proposed dual-pass GCN encoder improve node embeddings compared to standard GNN encoders? (6) hyperparameter sensitivity: How sensitive is GADL to variations in hyperparameter settings?
A comprehensive description of the experimental setup—including benchmarks, baselines, evaluation metrics, and implementation details—is provided in the Appendix C.
@@ 4.1 Robustness: evaluation on semi-synthetic benchmarks
To evaluate the robustness of the proposed GADL model under varying degrees of structural inconsistencies, we conduct experiments on six semi-synthetic graph alignment benchmark datasets: Celegans, Arena, Douban, Cora, DBLP, and Coauthor CS.
We follow the experimental setting introduced in the (He et al., 2024) for generating inconsistent graph pairs. Given a source graph $\mathcal{G}_{s}$ with adjacency matrix $\mathbf{A}$, we construct 10 perturbed and permuted target graphs using the transformation $\hat{\mathbf{A}}=\mathbf{P}(\mathbf{A}+\mathbf{M})\mathbf{P}^{\top}$, where $\mathbf{M}\in\{-1,0,1\}^{N\times N}$ introduces edge-level perturbations, and $\mathbf{P}$ is a random permutation matrix. The perturbation level is controlled by a parameter $p\in\{0,1\%,5\%\}$, representing the fraction of edges modified: $p|\mathcal{E}|$. We adopt seven structural node features from (Berlingerio et al., 2013): node degree, clustering coefficient, average degree of neighbors, average clustering coefficient of neighbors, number of edges in ego-network, number of outgoing edges of ego-network, and number of neighbors of ego-network. These descriptors provide compact, structure-aware representations for robust evaluation under varying structural inconsistencies.
Table 1 presents a comprehensive comparison of our proposed method, GADL, against six state-of-the-art network alignment methods: Netsimile (Berlingerio et al., 2013), Final (Zhang and Tong, 2016), GAlign (Trung et al., 2020), WAlign (Gao et al., 2021b), GAE (Kipf and Welling, 2016), T-GAE(He et al., 2024), and SLOTAlign (Tang et al., 2023a). Results show mean matching accuracy and standard deviation across 10 randomly generated target graphs under structural perturbations of 0%, 1%, and 5%. Results for Final, WAlign, and GAE are from He et al. (2024), while GAlign and SLOTAlign results are reproduced from their original implementations. Entries marked "–" indicate methods that failed to complete within reasonable time limits due to scalability constraints.
@@
The experimental results yield several key observations: (1) GADL consistently ranks among the top performers across all datasets, demonstrating robust performance under varying structural inconsistencies. While baseline methods often exhibit sharp performance degradation under structural noise, GADL maintains high accuracy even with 5% edge perturbations. (2) Embedding-based GA methods (e.g., GAE, T-GAE, GAlign, GADL) generally outperform optimal-transport-based methods (e.g., Final). SLOTAlign, which combines both learning-based and optimization techniques, achieves competitive but suboptimal results compared to pure learning-based models.
@@ 4.2 Effectiveness: evaluation on real-world benchmarks
We evaluate the effectiveness of the proposed GADL method on two real-world noisy graph datasets with partial node alignment: Douban Online-Offline and ACM-DBLP. These benchmarks reflect realistic and challenging scenarios, involving distinct graphs with partially aligned nodes. We evaluate the performance of GADL and baseline methods using Hit@$k$, which measures the proportion of ground-truth aligned nodes ranked among the top-$k$ predictions. The results are reported in Table 2.
@@
The results reveal the following key insights: (1) GADL consistently achieves the highest alignment accuracy across both real-world datasets, outperforming all baselines. In particular, it surpasses the second-best models—T-GAE on ACM-DBLP and SLOTAlign on Douban—by substantial margins, demonstrating its robustness in unsupervised real-world graph alignment scenarios. (2) Compared to T-GAE, which employs a GAE framework with a GIN encoder but lacks latent-space communication, our GADL model demonstrates superior performance through its dual-pass GCN encoder architecture integrated with latent-space communication. This design enables GADL to generate superior node embeddings, improving performance on unsupervised graph alignment tasks. (3) Beyond GADL, other learning-based methods (e.g., T-GAE) also outperform optimal-transport approaches (e.g., SLOTAlign) on larger benchmarks (e.g., ACM-DBLP), demonstrating superior robustness of learning-based methods to structural variations in large graphs that violate optimal transport assumptions.
@@ 4.3 Generalization: evaluation on vision–language benchmarks
Latent space alignment is a special case of graph alignment, relying only on embeddings without explicit structure. To highlight this generality, we further evaluate our method on vision-language alignment benchmarks, where the task involves aligning latent representations from diverse pretrained vision and language models.
We evaluate latent space alignment across multiple benchmarks using representations from diverse pretrained vision and language models. Full experimental details are provided in Appendix C.4.
Table 3 summarizes the vision-language alignment accuracies on four datasets using three pretrained vision models (CLIP (Ramesh et al., 2022), DeiT (Touvron et al., 2021), and DINOv2 (Oquab et al., 2023)) and two pretrained language models from SentenceTransformers library (Reimers and Gurevych, 2019) (all-mpnet-base-v2 and all-roberta-large-v1). Comprehensive results for all vision and language model combinations are provided in the Appendix F. Results on CIFAR-100 and ImageNet-100 are reproduced using official implementations.
@@
The experimental results highlight several key insights. (1) GADL consistently outperforms all baselines across datasets and pretrained models, demonstrating that moving beyond optimization- and optimal transport–based frameworks provides substantial benefits. (2) Most baselines achieve near-chance accuracies ($\leq 10\%$) with occasional inconsistent successes, and even sophisticated solvers like Gurobi or Hahn-Grant fail in certain settings. This inconsistency intensifies with dataset complexity, and the performance gap with GADL becomes pronounced on challenging benchmarks (CIFAR-100, ImageNet-100), underscoring the limitations of treating vision-language alignment as pure assignment optimization. (3) The choice of pretrained model critically impacts cross-modal alignment performance. While DINOv2 and DeiT achieve highest accuracies on smaller datasets, CLIP consistently outperforms on larger, more complex benchmarks. This highlights how pretrained model selection significantly influences visual-language alignment effectiveness.
4.4 Ablation analysis
To better understand the contribution of each component in the proposed GADL framework, we conduct an ablation study evaluating variants with specific modules removed or modified. We compare GADL against: (1) GADL w/o dual-pass encoder: replaces the dual-pass GCN with a standard single-pass GCN while retaining latent-space communication; (2) GADL w/o bijectivity regularization: removes the bijectivity regularization term while keeping the dual-pass encoder; (3) GADL w/o orthogonality regularization: omits orthogonality regularization while maintaining all other components.
Method ACM-DBLP Douban Online-Offline
Hit@1 Hit@5 Hit@10 Hit@50 Hit@1 Hit@5 Hit@10 Hit@50
GADL 88.63 94.76 96.16 98.41 53.31 73.61 80.67 94.18
The ablation results are summarized in Table 4. They highlight the individual contribution of each component to the overall alignment performance. Replacing the dual-pass GCN with a standard encoder causes substantial accuracy drops. Furthermore, the removal of the regularization terms notably degrades performance, confirming their importance in achieving robust and reliable node alignment.
Further analysis and experiments can be found in the supplementary material (Appendices D and E).
@@ 5 Conclusion
We present GADL, a novel framework for unsupervised graph alignment that combines dual-pass encoding with geometry-aware latent space communication. Comprehensive experiments demonstrate consistent performance gains across diverse benchmarks, with successful application to vision-language tasks validating the framework’s broader utility. While promising, our approach incurs modest computational overhead from dual-pass encoding compared to standard GCN and requires careful hyperparameter tuning. Future work will focus on adaptive spectral filtering and efficient embedding strategies, with potential extensions to molecular networks, social graphs, and multi-modal alignment tasks.
Appendices
@@ Appendix A Related work
Extensive research has addressed the graph alignment problem, with existing methods broadly categorized into three families based on their alignment strategies: optimization-based, optimal transport-based, and embedding-based approaches. These methods also differ in terms of the level of supervision required—ranging from unsupervised techniques to semi-supervised methods (which rely on partially paired nodes), and fully supervised approaches.
Traditional graph alignment methods formulate the problem as an optimization task, typically as a Quadratic Assignment Problem (QAP), seeking node permutations that minimize discrepancies between source and target adjacency matrices. IsoRank [Singh et al., 2008] represents a seminal approach, employing a PageRank-inspired algorithm to compute node similarity matrices based on neighbor similarity for unsupervised alignment. BigAlign [Koutra et al., 2013] extends this framework by incorporating both structural and attribute information to enhance alignment accuracy. FINAL [Zhang and Tong, 2016] addresses scalability through matrix factorization, combining global structural consistency with partial anchor constraints.
@@
Optimal transport-based methods model each graph as a probability distribution over its nodes and seek a transport plan—i.e., a soft correspondence—that minimizes a divergence such as the Wasserstein or Gromov-Wasserstein distance between the distributions. This framework offers a principled approach to graph alignment by optimizing the transport cost between node distributions. Unlike the hard alignments produced by QAP-based methods, optimal transport typically yields soft alignment matrices, allowing for uncertainty and partial correspondences.
A notable early contribution in this category is WAlign [Gao et al., 2021b], which jointly learns node embeddings and alignments by minimizing Wasserstein distance between graphs in a shared embedding space using a lightweight GCN and Wasserstein distance discriminator. Building on this direction, FGW [Tang et al., 2023b] employs Fused Gromov-Wasserstein distance to jointly align structural and attribute information through a coarse-to-fine matching scheme. PARROT [Zeng et al., 2023] extends this idea by running Random Walk with Restart (RWR) on both individual graphs and their Cartesian product, capturing more nuanced structural correspondence. GALOPA [Wang et al., 2023] integrates a GNN encoder with a self-supervised OT loss, jointly learning features and transport plans. To improve scalability, Wasserstein Wormhole [Haviv et al., 2024] introduces a transformer-based autoencoder that maps distributions into a latent space where Euclidean distances approximate Wasserstein distances, enabling efficient, linear-time graph comparisons.
@@
Embedding-based methods learn vector representations for nodes in each graph and align them based on embedding similarity. This approach typically involves generating node embeddings—either independently or jointly—followed by alignment through nearest-neighbor search or learned mapping functions.
NetSimile [Berlingerio et al., 2013] represents an early embedding-based approach that uses handcrafted structural features (degree, clustering coefficient) to represent nodes and aligns graphs through direct feature vector comparison using similarity measures. GAlign [Trung et al., 2020] adopts an unsupervised approach where both graphs are independently encoded using a shared Graph Convolutional Network, with node embeddings aligned by minimizing distributional discrepancies such as Wasserstein distance between embedding spaces. NeXtAlign [Zhang et al., 2021] enhances representation learning through a cross-graph attention mechanism that enables nodes in one graph to attend to features in the other. This produces alignment-aware embeddings and improves performance in semi-supervised settings with known anchor node pairs. REGAL [Heimann et al., 2018] generates compact node embeddings by extracting structural features like node degree and local neighborhoods, then aligns nodes across graphs by matching their embeddings based on distance, enabling efficient and scalable graph alignment. GINA [Wang et al., 2022b] addresses hierarchical alignment by projecting node embeddings from Euclidean to hyperbolic space, learning linear transformations between geometries using anchor nodes to better capture scale-free and hierarchical structures in social and biological networks.
A foundational approach to embedding-based graph learning is the Graph Autoencoder (GAE) and its probabilistic extension, the Variational Graph Autoencoder (VGAE) [Kipf and Welling, 2016]. These models use a GCN encoder to generate latent node embeddings, which are then used to reconstruct the adjacency matrix via an inner product decoder. Although originally designed for link prediction, GAEs have become a common backbone for alignment tasks due to their ability to capture global graph structure in an unsupervised manner. Expanding on this foundation, T-GAE [He et al., 2024] addresses scalability through a transferable graph autoencoder trained on small graph families that generalizes to large, unseen networks without fine-tuning. This design enables strong alignment performance while significantly reducing training time and computational overhead. However, typical embedding-based methods often become unstable when graphs differ significantly in structure. SLOTAlign [Tang et al., 2023a] is developed to tackle the structure and feature inconsistencies commonly found in these embedding-based graph alignment methods. It formulates alignment as an optimal transport problem on learned intra-graph similarity matrices, combining optimal transport with embedding-based approaches.

### v2

@@ 3.3 Latent space communication
While each GAE independently produces a latent space for its respective graph, resulting in misaligned embeddings, we address this limitation by incorporating deep functional maps to learn explicit mappings between latent representations. Rather than directly comparing raw embeddings, which may differ by arbitrary isometric transformations, we learn functional maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ that transform functions between latent spaces. These maps are optimized within our network using Equation 2, where $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ represent embeddings from the shared dual-pass encoder.
To facilitate latent space communication, our framework leverages spectral geometry principles and a regularized functional map module that enforces structural constraints. We impose bijectivity and orthogonality losses to ensure the maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ are approximately invertible and locally isometric, preserving essential geometric properties. The bijectivity loss promotes invertibility by ensuring functions mapped between latent spaces and back are accurately reconstructed, enforcing structural consistency and mutual alignment. Formally, it is defined as:
$\mathcal{L}_{\mathrm{bij}}=\|\mathbf{C}_{12}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}\mathbf{C}_{12}-\mathbf{I}\|_{F}^{2}.$ (4)
@@
$\mathcal{L}_{\mathrm{orth}}=\|\mathbf{C}_{12}\mathbf{C}_{12}^{\top}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}^{\top}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}.$ (5)
@@ 3.5 Model optimization and training Loss
Our model jointly optimizes GAE parameters and functional maps $\mathbf{C}_{12}$, $\mathbf{C}_{21}$ through end-to-end training. Given embeddings $\mathbf{Z}_{1}=f_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})$ and $\mathbf{Z}_{2}=f_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})$, we project them into spectral domains using graph Laplacian eigenvectors, yielding descriptors $\hat{\mathbf{F}}_{1}$ and $\hat{\mathbf{F}}_{2}$. Functional maps $\mathbf{C}_{12}\in\mathbb{R}^{k\times k}$ and $\mathbf{C}_{21}\in\mathbb{R}^{k\times k}$ align these spectral features via:
$\mathcal{L}_{\mathrm{FM}}^{12}=\alpha\left\|\mathbf{C}_{12}\hat{\mathbf{F}}_{1}-\hat{\mathbf{F}}_{2}\right\|_{F}^{2}+\beta\left\|\Lambda_{2}\mathbf{C}_{12}-\mathbf{C}_{12}\Lambda_{1}\right\|_{F}^{2}$ (6)
$\mathcal{L}_{\mathrm{FM}}^{21}=\alpha\left\|\mathbf{C}_{21}\hat{\mathbf{F}}_{2}-\hat{\mathbf{F}}_{1}\right\|_{F}^{2}+\beta\left\|\Lambda_{1}\mathbf{C}_{21}-\mathbf{C}_{21}\Lambda_{2}\right\|_{F}^{2}$ (7)
We incorporate these objectives as differentiable loss terms, with $\mathbf{C}_{12}$ and $\mathbf{C}_{21}$ as trainable parameters optimized end-to-end via backpropagation. These losses are combined with the standard GAE reconstruction loss, minimizing binary cross-entropy between the observed adjacency matrix $\mathbf{A}$ and its reconstruction $\hat{\mathbf{A}}$:
$\mathcal{L}_{\mathrm{rec}}=\mathrm{BCE}(\mathbf{A}_{s},\hat{\mathbf{A}}_{s})+\mathrm{BCE}(\mathbf{A}_{t},\hat{\mathbf{A}}_{t}),$ (8)
@@
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{rec}}+\lambda_{\mathrm{FM}}\left(\mathcal{L}_{\mathrm{FM}}^{12}+\mathcal{L}_{\mathrm{FM}}^{21}\right)+\lambda_{\mathrm{bij}}\,\mathcal{L}_{\mathrm{bij}}+\lambda_{\mathrm{orth}}\,\mathcal{L}_{\mathrm{orth}}$ (9)
@@ 4 Experiments
In this section, we aim to address the following research questions: (1) robustness: is GADL more robust to feature and structural inconsistencies than existing state-of-the-art graph alignment methods? (2) effectiveness: does GADL outperform state-of-the-art methods on real-world graph alignment tasks? (3) generalization: how effectively does GADL generalize to vision–language alignment? (4) ablation analysis: what is the contribution of each component in GADL to the overall alignment performance? (5) encoder evaluation: how does the proposed dual-pass GCN encoder improve node embeddings over standard GNNs? (6) hyperparameter sensitivity: how sensitive is GADL to hyperparameter variations?
A comprehensive description of the experimental setup, including benchmarks, baselines, evaluation metrics, and experimental settings, is provided in the Appendix C.
@@ 4.1 Robustness: evaluation on semi-synthetic benchmarks
To evaluate the robustness of the proposed GADL model under structural inconsistencies, we conduct experiments on six semi-synthetic benchmark datasets following (He et al., 2024), generating perturbed graph pairs with perturbation levels of 0%, 1%, and 5% (setup in Appendix C.3).
Table 1 compares GADL against state-of-the-arts: Netsimile (Berlingerio et al., 2013), Final (Zhang & Tong, 2016), GAlign (Trung et al., 2020), WAlign (Gao et al., 2021a), GAE (Kipf & Welling, 2016), T-GAE(He et al., 2024), and SLOTAlign (Tang et al., 2023a). Results show mean matching accuracy and standard deviation across 10 randomly generated target graphs under structural perturbations of 0%, 1%, and 5%. Results for Final, WAlign, and GAE are from He et al. (2024), while GAlign and SLOTAlign are reproduced. Entries marked ”–” indicate scalability failures.
@@
The results yield several key observations: (1) GADL consistently ranks among the top performers across datasets, maintaining high accuracy even with 5% perturbations while baselines show sharp degradation under structural noise. (2) Embedding-based methods (GAE, T-GAE, GAlign, GADL) generally outperform optimal-transport-based methods (Final). SLOTAlign, combining learning and optimization, achieves competitive but suboptimal results compared to pure learning-based models.
@@ 4.2 Effectiveness: evaluation on real-world benchmarks
We evaluate the effectiveness of the proposed GADL method on two real-world noisy graph datasets with partial node alignment: Douban Online-Offline and ACM-DBLP. These benchmarks involve distinct graphs with partially aligned nodes. Performance is measured using Hit@$k$, the proportion of ground-truth nodes ranked in the top-$k$ predictions. Results are reported in Table 2.
@@
The results reveal key insights: (1) GADL consistently achieves the highest alignment accuracy, outperforming all baselines, outperforming all baselines with substantial margins over second-best models (T-GAE on ACM-DBLP, SLOTAlign on Douban). (2) Compared to T-GAE, which employs a GIN encoder but lacks latent-space communication, our GADL model demonstrates superior performance through its dual-pass GCN encoder architecture integrated with latent-space communication. (3) Learning-based methods (T-GAE, GADL) outperform optimal-transport approaches (SLOTAlign) on larger benchmarks, demonstrating better robustness to structural variations that violate optimal transport assumptions.
Additional experiments on ablation analysis, encoder evaluation, and hyperparameter sensitivity are provided in Appendices D, E, and F.
@@ 4.3 Generalization: evaluation on vision–language benchmarks
Latent space alignment is a special case of graph alignment, relying only on embeddings without explicit structure. To highlight this generality, we further evaluate our method on vision-language alignment benchmarks, where the task involves aligning latent representations from diverse pretrained vision and language models. We evaluate latent space alignment across multiple benchmarks using representations from diverse pretrained vision and language models. Full experimental details are provided in Appendix C.4.
Table 3 summarizes the vision-language alignment accuracies on four datasets using three pretrained vision models (CLIP (Ramesh et al., 2022), DeiT (Touvron et al., 2021), and DINOv2 (Oquab et al., 2023)) and two pretrained language models from SentenceTransformers library (Reimers & Gurevych, 2019) (all-mpnet-base-v2 and all-roberta-large-v1). Comprehensive results are in the Appendix G. Results on CIFAR-100 and ImageNet-100 are reproduced using official implementations.
@@
The results highlight key insights. (1) GADL consistently outperforms all baselines, demonstrating substantial benefits beyond optimization and optimal transport frameworks. (2) Most baselines achieve near-chance accuracies ($\leq 10\%$) with occasional inconsistent successes, even sophisticated solvers like Gurobi fail in certain settings. Performance gaps with GADL become pronounced on challenging benchmarks (CIFAR-100, ImageNet-100), highlighting limitations of treating alignment as pure assignment optimization. (3) Pretrained model choice critically impacts performance, while DINOv2 and DeiT excel on smaller datasets, CLIP consistently outperforms on larger benchmarks.
@@ 5 Conclusion
We present GADL, a novel framework for unsupervised graph alignment that combines dual-pass encoding with geometry-aware latent space communication. Comprehensive experiments demonstrate consistent performance gains across diverse benchmarks, with successful application to vision-language tasks validating the broader utility of the framework beyond traditional graph domains. While promising, our approach incurs modest computational overhead from dual-pass encoding compared to standard GCN and requires careful hyperparameter tuning. Future work will focus on adaptive spectral filtering and efficient embedding strategies, with potential extensions to molecular networks, social graphs, and multi-modal alignment tasks, including a more thorough evaluation on vision–language benchmarks.
@@ Appendix A Related work
Extensive research has addressed the graph alignment problem, with existing methods broadly categorized into three families based on their alignment strategies: optimization-based, optimal transport-based, and embedding-based approaches. These methods also differ in terms of the level of supervision required, ranging from unsupervised techniques to semi-supervised methods (which rely on partially paired nodes), and fully supervised approaches.
Traditional graph alignment methods formulate the problem as an optimization task, typically as a Quadratic Assignment Problem (QAP), seeking node permutations that minimize discrepancies between source and target adjacency matrices. IsoRank (Singh et al., 2008) represents a seminal approach, employing a PageRank-inspired algorithm to compute node similarity matrices based on neighbor similarity for unsupervised alignment. BigAlign (Koutra et al., 2013) extends this framework by incorporating both structural and attribute information to enhance alignment accuracy. FINAL (Zhang & Tong, 2016) addresses scalability through matrix factorization, combining global structural consistency with partial anchor constraints.
@@
Optimal transport-based methods model each graph as a probability distribution over its nodes and seek a transport plan, i.e., a soft correspondence, that minimizes a divergence such as the Wasserstein or Gromov-Wasserstein distance between the distributions. This framework offers a principled approach to graph alignment by optimizing the transport cost between node distributions. Unlike the hard alignments produced by QAP-based methods, optimal transport typically yields soft alignment matrices, allowing for uncertainty and partial correspondences.
A notable early contribution in this category is WAlign (Gao et al., 2021a), which jointly learns node embeddings and alignments by minimizing Wasserstein distance between graphs in a shared embedding space using a lightweight GCN and Wasserstein distance discriminator. Building on this direction, FGW (Tang et al., 2023b) employs Fused Gromov-Wasserstein distance to jointly align structural and attribute information through a coarse-to-fine matching scheme. PARROT (Zeng et al., 2023) extends this idea by running Random Walk with Restart (RWR) on both individual graphs and their Cartesian product, capturing more nuanced structural correspondence. GALOPA (Wang et al., 2023) integrates a GNN encoder with a self-supervised OT loss, jointly learning features and transport plans. To improve scalability, Wasserstein Wormhole (Haviv et al., 2024) introduces a transformer-based autoencoder that maps distributions into a latent space where Euclidean distances approximate Wasserstein distances, enabling efficient, linear-time graph comparisons.
@@
Embedding-based methods learn vector representations for nodes in each graph and align them based on embedding similarity. This approach typically involves generating node embeddings, either independently or jointly, followed by alignment through nearest-neighbor search or learned mapping functions.
NetSimile (Berlingerio et al., 2013) represents an early embedding-based approach that uses handcrafted structural features (degree, clustering coefficient) to represent nodes and aligns graphs through direct feature vector comparison using similarity measures. GAlign (Trung et al., 2020) adopts an unsupervised approach where both graphs are independently encoded using a shared Graph Convolutional Network, with node embeddings aligned by minimizing distributional discrepancies such as Wasserstein distance between embedding spaces. NeXtAlign (Zhang et al., 2021) enhances representation learning through a cross-graph attention mechanism that enables nodes in one graph to attend to features in the other. This produces alignment-aware embeddings and improves performance in semi-supervised settings with known anchor node pairs. REGAL (Heimann et al., 2018) generates compact node embeddings by extracting structural features like node degree and local neighborhoods, then aligns nodes across graphs by matching their embeddings based on distance, enabling efficient and scalable graph alignment. GINA (Wang et al., 2022b) addresses hierarchical alignment by projecting node embeddings from Euclidean to hyperbolic space, learning linear transformations between geometries using anchor nodes to better capture scale-free and hierarchical structures in social and biological networks.
A foundational approach to embedding-based graph learning is the Graph Autoencoder (GAE) and its probabilistic extension, the Variational Graph Autoencoder (VGAE) (Kipf & Welling, 2016). These models use a GCN encoder to generate latent node embeddings, which are then used to reconstruct the adjacency matrix via an inner product decoder. Although originally designed for link prediction, GAEs have become a common backbone for alignment tasks due to their ability to capture global graph structure in an unsupervised manner. Expanding on this foundation, T-GAE (He et al., 2024) addresses scalability through a transferable graph autoencoder trained on small graph families that generalizes to large, unseen networks without fine-tuning. This design enables strong alignment performance while significantly reducing training time and computational overhead. However, typical embedding-based methods often become unstable when graphs differ significantly in structure. SLOTAlign (Tang et al., 2023a) is developed to tackle the structure and feature inconsistencies commonly found in these embedding-based graph alignment methods. It formulates alignment as an optimal transport problem on learned intra-graph similarity matrices, combining optimal transport with embedding-based approaches.